solve for t using natural logarithms.
4e^9t=7e^8t

Answer by lwsshak3(11628)

solve for t using natural logarithms.
4e^9t=7e^8t
ln4+9t*lne=ln7+8t*lne
lne=1 (log of base=1)
ln4+9t=ln7+8t
t=ln7-ln4
t≈0.5596